.600g of solid CO2 is put in an empty sealed 4.00L container at a temperature of 300K.When all the solid CO2 becomes gad, what will be the pressure in the container?


1
Expert's answer
2021-04-05T07:51:23-0400

no of moles of CO2 = .600/44.01 = 0.0136mol

PV = nRT

P × 4 = 0.0136×0.0821 × 300

P × 4 = 0.335

P = 0.08375 atm